OpenStudy (anonymous):

Convert the spherical polar equation p=4cosɸ to rectangular coordinates and identify the graph

OpenStudy (nikvist):

\[\rho=4\cos{\phi}\]\[\rho^2=4\rho\cos{\phi}\]\[x^2+y^2=4x\]\[(x-2)^2+y^2=2^2\]